PART A) A 0.20 kg mass on a horizontal spring is pulled back 2.0 cm and...

PART A)

A 0.20 kg mass on a horizontal spring is pulled back 2.0 cm and released. If, instead, a 0.40 kg mass were used in this same experiment, the total energy of the system would

A) Remain the same. B) Double. C) Be half as large.

PART B)

A 0.20 kg mass on a horizontal spring is pulled back a certain distance and released. The maximum speed of the mass is measured to be 0.30 m/s. If, instead, a 0.40 kg mass were used in this same experiment, choose the correct value for the maximum speed.

A) 0.21 m/s. B) 0.60 m/s. C) 0.42 m/s. D) 0.15 m/s. E) 0.30 m/s.
Answer #1

A) energy of the spring = 0.5kx^2

so it depends on spring constant

and change in length

so it is independent of mass

so A) remains the same is the correct answer

B) total energy = 0.5mv^2= 0.5*0.2*0.3^2

= 9*10^-3 J

9*10^-3 = 0.5*0.4*v^2

v = 0.21 m/s

so option A) 0.21 m/s is the correct answer